OpenStudy (anonymous):

Solve: x^4 - 6x^2 + 5 = 0

OpenStudy (anonymous):

fortunately this one factors as \[(x^2-5)(x^2-1)=0\]

OpenStudy (anonymous):

solve for \(x\) for example \[x^2-5=0\implies x^2=5\implies x=\pm\sqrt5\]

OpenStudy (anonymous):

So x would equal +/-1 & +/- \[\sqrt{5}\]?